A car of mass ma is driving around with speed v in a circle of radius R on a banked, frictionless road at an angle θ above the horizontal. 3.
Answer #1

for finding the angle of banking

tan \theta = v^2 / rg

where \theta = angle of banking

v = velocity at turn

r = radius , g = acceleration due to gravity